# Video In-context Learning

### v1

@@ 4 Experimental Setups
We design tailored evaluation pipelines to assess the model capacity in Vid-ICL. We introduce the proposed evaluation metrics in Section 4.1, and summarize the details of our implementations in Section 4.2.
@@ 4.1 Evaluation
The evaluation of Vid-ICL should include two aspects, the first is the visual quality as a normal video generation task, and the second is semantic accuracy which is more critical as it reflects whether the model understands and follows the semantic guidance provided by demonstrations. To validate semantic accuracy, we design four kinds of demonstrations.
  • No Demonstration. This setting is akin to normal video continuation, where the model is asked to complete the query $x^{V}$ without any in-context demonstration.
@@
For visual quality, we adopt several widely utilized metrics including PSNR and LPIPS [46] to perform a frame-by-frame comparison between the generation results and the ground truth video clip, i.e., the consequences of the query. Additionally, we include the FID score [16] to quantify the distribution difference. These metrics are suitable for the first setting where ground truth is available.
@@
  • Video Accuracy (V-Acc). We utilize an off-the-shelf video classifier [33] trained on the SSv2 dataset to calculate the classification accuracy of the generated video clips. Specifically, we predict the action label of each generated video clip and compare it with the ground truth label of the query. This metric provides a perceptual assessment of the semantic information in generation results.
  • Probing Accuracy (P-Acc). V-Acc relies on a pretrained classifier and judges on visual signals. To directly validate the semantic information contained in latent representations, inspired by the widespread use of probing in vision feature extractors [15, 22, 4], we train a probing classifier which takes hiddens in the last Transformer layer of the generation results as input and predicts the corresponding action label.
@@ 4.2 Implementation Details
For each dataset, we assign a stride to sample frames at regular intervals in order to capture human-recognizable video clips. The stride varies among datasets depending on the average FPS but remains consistent within each dataset. Subsequently, we resize and center-crop the images to a fixed size of $256\times 256$ resolution.The pretrained VQ-GAN tokenizer [28] takes in $256\times 256$ sized images and produces $16\times 16=256$ discrete tokens with a compression coefficient $f=16$. Tokens of different frames are concatenated in the order of original frames, and each training sequence contains 16 images and 4096 discrete tokens.
@@
we adopt the LlaMA [34] architecture, a typical decoder-only transformer model for auto-regressive modeling with designs advantageous to large-scale modeling. we experiment with different sets of hyper-parameters which result in models with 300M, 700M and 1.1B parameters separately. The details of hyper-parameter settings are shown in Table 9.
@@
LVM [1] is a natural baseline that shows strong Img-ICL capacity. However, due to the absence of an open-source checkpoint and the high resource demands required to reproduce the model, we utilize an open reproduction, DeLVM [13], instead. To evaluate the generalizability of our model, we train several variants: Pretrain, which is trained without SSv2 in the training set; Pretrain w/ In-domain finetune, which is Pretrain fine-tuned on SSv2 where each sample consists of continuous 16 frames, similar to the pretraining stage; and Pretrain w/ In-context finetune, which is fine-tuned using the same data format as in-context inference, consisting of two 8-frame videos from the same class.
@@ 5 Results and Analysis
V-Acc $\uparrow$ P-Acc $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
@@
Random 22.7 28.3 13.01 0.450 22.53
No 22.9 29.6 13.20 0.442 20.94
In-class 24.7 36.7 13.07 0.447 21.95
Pretrain w/ In-context Finetune
Random 23.1 25.6 13.01 0.454 20.66
No 24.2 26.7 13.02 0.460 22.21
In-class 25.7 40.7 13.08 0.450 20.49
@@
Random 24.9 34.1 13.16 0.440 19.17
No 25.0 30.8 13.09 0.446 20.87
In-class 25.9 48.8 13.21 0.438 18.92
@@ 5.1 Main results
If not specified, we use our largest 1.1B model to generate results. We first evaluate the Vid-ICL results using the ground truth query label. In this setting, all automatic metrics can be utilized. The quantitative results are presented in Table 1, where each row corresponds to a different demonstration type. We draw the following conclusions from the results.
In-class demonstrations contribute to stronger semantic accuracy Our analysis begins by focusing on the result of the Pretrain model in the first block. We observe a significant enhancement in the semantic accuracy when using in-class video clips as demonstrations, compared to without demonstration or using randomly selected ones. Specifically, there is a notable 7.1% / 8.4% improvement in P-Acc and a 1.8% / 2.0% gain in V-Acc over no or prompting with random demonstrations respectively. These findings validate our proposition that when prompted with semantically related demonstrations, the model more accurately generates video clips that adhere to the original trace.
@@
In-domain finetuning leads to general improvement We further analyze the results presented in the second and third blocks of Table 1, focusing on different fine-tuning strategies. The superiority of in-domain finetuning over the initial Pretrain is evident, indicating a straightforward performance boost when in-domain training data is available. In addition, when comparing in-domain finetuning with in-context finetuning, the in-class P-Acc of the former rises to 48.8%, showcasing an 8.1% improvement over the latter. These findings underscore the remarkable zero-shot capacity of Vid-ICL, showcasing the model’s capability at adapting to in-context formatting through exclusive training on continuous video clips, while explicit in-context finetuning fails to yield significant enhancements.
@@
Evaluation with demonstration label We evaluate the generation results with the demonstration label to assess the controliability of the model. We use results with both in-class and contrastive demonstrations as described in Section 4.1 to train the probing model. As shown in Table 2, both cases have well above average probing accuracy, indicating that the representations of generation results successfully obtain the semantic information from demonstrations.
Table 3: Results on automatic and human evaluation of Vid-ICL and DeLVM baseline.
P-Acc PSNR Quality Alignment Control
DeLVM 32.8 12.69 3.21 2.28 2.11
Vid-ICL 38.5 13.07 4.12 3.73 3.01
5.2 Comparison with DeLVM Baseline
DeLVM [13] accepts a maximum of $8$ frames in the context window. To align with this, we adapt our model to receive 4 frames of demonstration, 2 frames of query, and generate 2 frames as the result. Both models contain approximately 1.1B parameters. We report both automatic and human evaluation results in Table 3.
As shown in the results, our Vid-ICL model significantly outperforms DeLVM in both objective and subjective evaluations. We attribute this difference to the training data formats of DeLVM (i.e., Img-ICL) and Vid-ICL. Img-ICL relies on pairs of images and their annotations, which limits its ability to extract semantic information across consecutive frames and generate coherent sequences correspondingly.
@@ 5.4 Ablation and Analysis
We investigate the impact of different numbers of demonstrations on Vid-ICL performance. Within a 16-frame context window, we fix the query length as 4 frames and the generation as 4 frames, and construct four demonstration formats: 1) no demonstration; 2) one 4-frame demonstration; 3) two 4-frame demonstrations sampled from different videos; 4) one 8-frame demonstration. The results in Table 5 demonstrate that providing demonstrations generally improves both P-Acc and V-acc scores. Additionally, providing more ($2\times 4$ versus $1\times 4$) or longer demonstrations ($1\times 8$ versus $1\times 4$) both result in better performance.
@@
To demonstrate the versatility of our model, we explore incorporating text into the demonstrations. We fine-tune the 300M Pretrain model on the SSv2 dataset by appending corresponding text annotations to the front of each video. The text annotations are encoded into latent representations using a pretrained T5-large model [30]. The results, presented in Table 5, indicate that our model can seamlessly adopt demonstrations in various modalities. Additionally, adding textual descriptions enhances the model’s in-context learning capacity.
Figure 4: The validation perplexity in various amount of web videos.
Effect of Web Videos
In previous studies, incorporating miscellaneous web videos into the training dataset has been shown to enhance generation quality [41, 26]. Here we investigate their effect on Vid-ICL. We sampled varying quantities of video clips from the Webvid dataset during pretraining and present the SSv2 validation perplexity results in Figure 4. We observe that the perplexity does not linearly decrease with the number of web videos, suggesting that Vid-ICL prefers videos conveying clear and consistent semantic information.
@@ 6 Conclusion
In this paper, we introduce Video In-Context Learning (Vid-ICL), a novel framework that extends in-context learning to video data. By converting video frames into sequences of discrete tokens and training an autoregressive Transformer with next token prediction, our model exhibits zero-shot in-context learning capabilities, which can generate semantically coherent and contextually appropriate video sequences based on provided demonstrations. Extensive experiments confirm that Vid-ICL effectively captures and conveys the semantic information embedded in the demonstrations, demonstrating its potential to enhance various downstream tasks such as visual planning. Additionally, the model’s versatility is verified by its ability to understand and integrate multi-modal demonstrations simultaneously.
Limitations Despite the promising results, several limitations exist in this study. Firstly, while Vid-ICL excels at generating short video sequences, its performance on longer sequences has not been thoroughly evaluated. Secondly, the model’s dependence on the quality and relevance of provided demonstrations can lead to inconsistencies, particularly when demonstrations are noisy or semantically misaligned. Future work should address these limitations by utilizing visual tokenizers with temporal compression, and scaling or developing robust models to handle imperfect demonstrations.
@@ Appendix A Additional Results
In this section we provide more samples on different video datasets and benchmarks that Vid-ICL model produces. Synthesis videos on Something-something v2 dataset are presented in Table 6. Table 7 includes samples on Robotics Transformer-1 dataset, showcasing the ability of Vid-ICL to fit on various form of video datasets. Furthermore, Vid-ICL as a versatile generalist, can act as a simulator in reinforcement learning tasks, for which we provide detailed explanations.
Vid-ICL as Simulator
@@
We demonstrate that Vid-ICL can also function as a simulator in reinforcement learning tasks by evaluating it on the VP2[32] benchmark. VP2 is a benchmark for video prediction models used in robotic manipulation via model-predictive control. Vid-ICL generates future frames with videos that accomplishing the same task as the demonstration, where the generated frames inversely reflect the corresponding actions that correctly interact with the environment. We compare the trajectories produced by the SVG baseline and Vid-ICL in Figure5. Evaluating on the Push-red task, we find that Vid-ICL provides more precise control over the environment interaction.
Figure 5: The trajectory produced by Svg and Vid-ICL respectively.
Table 6: More examples of Vid-ICL generated samples on Something-something v2 dataset.
Demonstration Query & Generation
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Table 7: More examples of Vid-ICL generated samples on RT-1 dataset.
Demonstration Query & Generation
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]

### v2

@@ 4 Experimental Setups
We design tailored evaluation pipelines to assess the VidIT’s capacity to perform video imitation. We introduce the proposed evaluation metrics in Section 4.1, and summarize the details of our implementations in Section 4.2.
@@ 4.1 Evaluation
The evaluation of VidIT should include two aspects, the first is the visual quality as a normal video generation task, and the second is semantic accuracy which is more critical as it reflects whether the model understands and follows the semantic guidance provided by demonstrations. To validate semantic accuracy, we design four kinds of demonstrations.
  • No Demonstration. This setting is akin to unconditional video prediction, where the model is asked to complete the query $x^{V}$ without any demonstration.
@@
For visual quality, we adopt several widely utilized metrics including PSNR and LPIPS (Zhang et al., 2018) to perform a frame-by-frame comparison between the generation results and the ground truth video clip, i.e., the consequences of the query. Additionally, we include the FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2018) score to quantify the distribution difference. These metrics are suitable for the first setting where ground truth is available.
@@
  • Video Accuracy (V-Acc). We utilize an off-the-shelf video classifier (Tong et al., 2022) trained on the SSv2 dataset to calculate the classification accuracy of the generated video clips. Specifically, we predict the action label of each generated video clip and compare it with the ground truth label of the query. This metric provides a perceptual assessment of the semantic information in generation results. In particular, we use the videomae-base-finetuned-kinetics [^2] https://huggingface.co/MCG-NJU/videomae-base-finetuned-kinetics checkpoint.
  • Probing Accuracy (P-Acc). While V-Acc relies on a pretrained classifier and judges on visual signals, we propose another metric named P-Acc to operate on the latent representation of VidIT. To directly validate the semantic information contained in latent representations, inspired by the widespread use of probing in vision feature extractors (He et al., 2022; Huang et al., 2023; Bardes et al., 2023), we train a probing classifier which takes hiddens in the last Transformer layer of the generation results as input and predicts the corresponding action label.
@@ 4.2 Implementation Details
For each dataset, we assign a stride to sample frames at regular intervals in order to capture human-recognizable video clips. The stride varies among datasets depending on the average FPS but remains consistent within each dataset. Subsequently, we resize and center-crop the images to a fixed size of $256\times 256$ resolution.The pretrained VQ-GAN tokenizer (Patil et al., 2024) takes in $256\times 256$ sized images and produces $16\times 16=256$ discrete tokens with a compression coefficient $f=16$. Tokens of different frames are concatenated in the order of original frames, and each training sequence contains 16 images and 4096 discrete tokens.
@@
we adopt the LLaMA (Touvron et al., 2023) architecture, a typical decoder-only transformer model for auto-regressive modeling with designs advantageous to large-scale modeling. we experiment with different sets of hyper-parameters which result in models with 300M, 700M and 1.1B parameters separately. The details of hyper-parameter settings are shown in Table 8. We introduce more training details in Appendix B.4
@@
We compare the proposed VidIT model with recent visual generative models based on auto-regressive Transformer. LVM (Bai et al., 2023) is a natural baseline that shows strong image imitation capacity. DeLVM (Guo et al., 2024) is a data efficient version of LVM, which involves extra knowledge distilling step and have fewer parameters. LWM (Liu et al., 2024) is designed to understand and generate long-context video contents, which is trained on aligned text-video pairs. To evaluate the generalizability of our model, we train several variants: Pretrain, which is trained without SSv2 in the training set; Pretrain w/ In-domain finetune, which is fine-tuned on SSv2 from the Pretrain model where each sample consists of continuous 16 frames, similar to the pretraining stage; and Pretrain w/ Imitation finetune, which is fine-tuned using the same data format as imitation inference, consisting of two 8-frame videos from the same class.
@@ 5 Results and Analysis
V-Acc $\uparrow$ P-Acc $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ FVD $\downarrow$
@@
No Demonstration 22.9 29.6 13.20 0.442 20.94 119.51
Random Demonstration 22.7 28.3 13.01 0.450 22.53 131.34
In-class Demonstration 24.7 36.7 13.07 0.447 21.95 125.77
Pretrain w/ Imitation Finetune
No Demonstration 24.2 26.7 13.02 0.460 22.21 129.68
Random Demonstration 23.1 25.6 13.01 0.454 20.66 115.76
In-class Demonstration 25.7 40.7 13.08 0.450 20.49 113.52
@@
No Demonstration 25.0 30.8 13.09 0.446 20.87 108.12
Random Demonstration 24.9 34.1 13.16 0.440 19.17 95.72
In-class Demonstration 25.9 48.8 13.21 0.438 18.92 88.63
@@ 5.1 Main results
If not specified, we use our largest 1.1B model to generate results. We first evaluate the VidIT results using the ground truth query label. In this setting, all automatic metrics can be utilized. The quantitative results are presented in Table 1, where each row corresponds to a different demonstration type. We draw the following conclusions from the results.
In-class Demonstrations Contribute to Stronger Semantic Accuracy Our analysis begins by focusing on the result of the Pretrain model in the first block. We observe a significant enhancement in the semantic accuracy when using in-class video clips as demonstrations, compared to without demonstration or using randomly selected ones. Specifically, there is a notable 7.1% / 8.4% improvement in P-Acc and a 1.8% / 2.0% gain in V-Acc over no or prompting with random demonstrations respectively. These findings validate our proposition that when prompted with semantically related demonstrations, the model more accurately generates video clips that adhere to the original trace, in a zero-shot way on both the domain and the imitation ability.
@@
In-domain Finetuning Leads to General Improvement We further analyze the results presented in the second and third blocks of Table 1, focusing on different fine-tuning strategies. The superiority of in-domain finetuning over the initial Pretrain is evident, indicating a straightforward performance boost when in-domain training data is available. In addition, when comparing in-domain finetuning with imitation finetuning, the in-class P-Acc of the former rises to 48.8%, showcasing an 8.1% improvement over the latter. These findings underscore the remarkable zero-shot capacity of VidIT, showcasing the model’s capability at adapting to the imitation formatting through exclusive training on continuous video clips, while explicit finetuning with the target format fails to yield significant enhancements.
Table 2: Results on automatic and human evaluation of VidIT and baseline models.
Baselines Automatic Metrics Human Evaluation
P-Acc PSNR Quality Alignment Control
LVM (Bai et al., 2023) 27.1 12.45 3.50 2.71 1.81
DeLVM (Guo et al., 2024) 32.8 12.69 3.21 2.28 2.11
LWM (Liu et al., 2024) 24.9 11.89 3.39 2.33 2.05
VidIT (ours) 38.5 13.07 4.12 3.73 3.01
@@
Evaluation with Demonstration Label We evaluate the generation results with the demonstration label to assess the controllability of the model. We use results with both in-class and contrastive demonstrations as described in Section 4.1 to train the probing model. As shown in Table 3, both cases have well above average probing accuracy, indicating that the representations of generation results successfully obtain the semantic information from demonstrations.
5.2 Comparison with Baseline Models
We compare our VidIT with baseline methods mentioned in Section 4.2 on the video imitation abilities.
We report both automatic metrics and human evaluation results in Table 2. The full evaluation of baseline models is reported in Table 6 in Appendix A.6. As shown in the results, our VidIT model significantly outperforms the baseline models in both objective and subjective evaluations. We attribute this difference to the training data formats of VidIT and other models. LVM and DeLVM mainly relies on pairs of images and their annotations, which limits their ability to extract semantic information across consecutive frames and generate coherent sequences correspondingly. Moreover, the training process of LWM are not designed to capture the demonstration’s semantics, leading to degradation on imitation performance.
@@ 5.3 Scaling Behavior
5.4 Generalization Ability
In this section, we elaborate on the generalization ability of the VidIT model. Thanks to its self-supervised training paradigm, VidIT can be smoothly adapted to multiple scenarios and tasks with minimal modifications.
Firstly, we apply VidIT to Robotics Transformer dataset (Brohan et al., 2022) to imitate actions of a robotic arm, such as opening and closing drawers and placing objects. The samples on Figure 4 demonstrates that VidIT can successfully mimic various robotic arm actions from demonstrations, completing the frame sequences with the same movements.
Additionally, we highlight that VidIT can serve as an underlying video generation engine within an interactive agent. In this setup, VidIT receives previous observations and generates subsequent frames, which are then converted to action signals via the inverse dynamics mechanism. This process ensures that the action signals reflect VidIT’s imitation intent. As depicted in Figure 6 in Appendix A.1, agents perform specific actions guided by VidIT, showcasing the zero-shot imitation ability of VidIT and revealing its potential to interact with the environment.
We also validate VidIT’s capability as a video task imitator. Following the approach of image task imitators like LVM, we create video segmentation exemplars to form QA pairs on the VIPSeg (Miao et al., 2022) dataset. The results, presented in Figure 7 in Appendix A.2, indicate that VidIT successfully learns the task format from the demonstrations. Moreover, we study how the model emerges the imitation ability by analyzing the attention score in Appendix A.5
Figure 4: Generation Samples of VidIT on RT-1 Dataset.
@@ 5.5 Ablation and Analysis
We investigate the impact of different numbers of demonstrations on VidIT performance. Within a 16-frame context window, we fix the query length as 4 frames and the generation as 4 frames, and construct four demonstration formats: 1) no demonstration; 2) one 4-frame demonstration; 3) two 4-frame demonstrations sampled from different videos; 4) one 8-frame demonstration. The results in Table 5 demonstrate that providing demonstrations generally improves both P-Acc and V-acc scores. Additionally, providing more ($2\times 4$ versus $1\times 4$) or longer demonstrations ($1\times 8$ versus $1\times 4$) both result in better performance.
@@
To demonstrate the versatility of our model, we explore incorporating text into the demonstrations. We fine-tune the 300M Pretrain model on the SSv2 dataset by appending corresponding text annotations to the front of each video. The text annotations are encoded into latent representations using a pretrained T5-large model (Raffel et al., 2020). The results, presented in Table 5, indicate that our model can seamlessly adopt demonstrations in various modalities. Additionally, adding textual descriptions enhances the model’s imitation capacity.
@@ 6 Conclusion
In this paper, we introduce Video Imitator, a novel model that extends imitation learning to video data. By converting video frames into sequences of discrete tokens and training the autoregressive Transformer with next token prediction, our VidIT model exhibits zero-shot imitation capabilities, which can generate semantically coherent and contextually appropriate video sequences based on provided demonstrations. Extensive experiments confirm that VidIT effectively captures and conveys the semantic information embedded in the demonstrations, demonstrating its potential to enhance various downstream tasks such as visual planning. Additionally, the model’s versatility is verified by its ability to understand and integrate multi-modal demonstrations simultaneously.
Acknowledgment
This research was supported by the National Natural Science Foundation of China (Grant No. 62276245).
@@ Appendix A Additional Results
In this section we provide more samples on different video datasets and benchmarks that VidIT model produces. Synthesis videos on Something-something v2 dataset are presented in Table 10. One can also navigate to our project page https://aka.ms/vid-icl for clearer presented samples. Table 11 includes samples on Robotics Transformer-1 dataset, showcasing the ability of VidIT to fit on various form of video datasets. Furthermore, VidIT as a versatile generalist, can act as a simulator in reinforcement learning tasks, for which we provide detailed explanations.
A.1 VidIT as Simulator
@@
We demonstrate that VidIT can also function as a simulator in reinforcement learning tasks by evaluating it on the VP2 (Tian et al., 2023) benchmark. VP2 is a benchmark for video prediction models used in robotic manipulation via model-predictive control. VidIT generates future frames with videos that accomplishing the same task as the demonstration, where the generated frames inversely reflect the corresponding actions that correctly interact with the environment (i.e. inverse dynamics). We compare the trajectories produced by the SVG (Villegas et al., 2019) baseline and VidIT in Figure 5. Evaluating on the Push-red task, we find that VidIT provides more precise control over the environment interaction.
Figure 5: The trajectory produced by SVG and VidIT respectively.
Additionally, we evaluate VidIT on the MineRL environment as the underlying video generation model, where VidIT take in previous game scenes as conditions to imitate and generate subsequent scenes. Generated game scenes are then fed into inverse dynamics models to generate the next action signal. We showcase two types of actions: Chop wood and Attack sheep. In Figure 6, the VidIT correctly generate game scenes and produce corresponding action signal to finish the action.
Figure 6: The VidIT model functions as video generator for chop-wood and attack-sheep in MineRL.
A.2 VidIT as Task Imitator
We expand the role of VidIT model to the task imitator on video, which mirrors the cases in image imitation learning. In this section, we finetune VidIT in VIPSeg dataset (Miao et al., 2022) for a few steps and show that VidIT can also smoothly generalize to video perceptual tasks like video segmentation. As depicted in Figure 7, VidIT successfully learns the task format from demonstration, and generates the segmentation outputs of the query video.
Figure 7: The VidIT can also imitate video perceptual tasks such as video segmentation.
A.3 Case Study: Baseline methods
In this section we compare the generated cases of VidIT with the baseline methods. As in Figure 8, we see that conditioned on the same frames, VidIT is capable of generating more semantically accurate and high-quality video than baseline models. Moreover, VidIT is lower in parameter budget, highlighting the priority of VidIT as a video imitator model.
Figure 8: Generated cases of VidIT and baseline methods.
A.4 Failure Cases
When prompted with semantically unrelated videos, VidIT may struggle to generate coherent content. In this section, we present failure cases where VidIT fails to align the query video with the semantic context of the demonstration video (see Figure 9).
For instance, in the first row, the demonstration video conveys the semantic information of moving a ball downward. However, when starting with a query video depicting an unrelated action, such as folding a blanket, the model struggles to reconcile the two contexts. As a result, it tends to ignore the demonstrated actions and generates content that lacks coherence with the demonstration.
Similar patterns are observed in other cases. These examples reveals VidIT’s potential difficulty in maintaining semantic coherence when prompted with unrelated or ambiguous demonstrations. Addressing this limitation presents a promising direction for future research.
Figure 9: Failure cases of VidIT.
Figure 10: The attention weight visualization during training and inference respectively.
A.5 Interpretability of Generalization
In this section, we highlight the interpretability of VidIT and its ability to generalize to imitation tasks in a zero-shot manner. To demonstrate this, we analyze the attention weights from the language model backbone to compare the similarity of attention patterns when provided with the same demonstration during training and inference.
Specifically, we extract the attention weights from the last layer of the model and average them across all attention heads. Since the tokens are flattened from the original frames in raster scan order, we rebind the tokens to their corresponding regions in the frames for better visualization. The results are shown in Figure 10.
The first row depicts the original frame sequence. The second row represents the training case, where frames are drawn from a single video clip. The third row illustrates the inference case, where the sequence begins from a different query frame. When generating a new token, we observe similar attention patterns across the demonstration frames. Notably, the attention is concentrated around the moving object within the frames.
This phenomenon indicates that the model effectively learns and leverages the semantic information from previous frames, enabling it to perform imitation tasks even when starting from a different scene.
Figure 11: The aggragated attention score distribution in demonstration and query frames.
Additionally, we analyze the failure cases in detail from the perspective of attention weights. Specifically, we examine how new frames attend to previous frames during generation. Intuitively, when a new frame is effectively guided by the demonstration, it tends to allocate more attention to the frames in the demonstration video, rather than only attending to the query frames.
To illustrate this, we visualize the attention distribution of generated frames on previous frames, as shown in Figure 11. Clearly, successful generation samples exhibit a tendency to allocate more attention to the demonstration frames, indicating an ability to comprehend and utilize its semantic information. In contrast, the attention curve for failure cases reveals a higher focus on the query frames, effectively ignoring the guidance provided by the demonstration part.
A.6 Full Evaluation Metrics on Baselines
In this section, we complete Table 2’s results with the all the evaluation metrics in Section 4.1. From the results in Table 6, it is evident that the VidIT model consistently outperforms the baseline models in video imitation tasks, reinforcing its superior capability in capturing both semantic alignment and visual quality.
Table 6: Full automatic metrics for baseline models. The best result is bolded .
Demonstration Semantic Accuracy Visual Quality
V-Acc $\uparrow$ P-Acc $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ FVD $\downarrow$
LVM (Bai et al., 2023) 21.2 27.1 12.45 0.489 27.58 163.13
DeLVM (Guo et al., 2024) 24.1 32.8 12.69 0.462 23.65 132.09
LWM (Liu et al., 2024) 21.8 24.9 11.89 0.474 29.91 191.56
VidIT (ours) 25.9 38.5 13.07 0.450 22.30 126.32
Appendix B Implementation Details